OpenStudy (anonymous):

Define \(f:[0,1]\to \mathbb{R}\) such that \(f\) is only continuous at one point. Furthermore, prove that the function you have provided is only continuous at that single point.

OpenStudy (anonymous):

I will provide a hint, and later a solution if no one can find one.

myininaya (myininaya):

\[f(x)=x, [-1,-1] \] lol

OpenStudy (anonymous):

Sorry, I don't get it

OpenStudy (anonymous):

That's not in the interval he gave.

myininaya (myininaya):

oh oops

OpenStudy (anonymous):

Let me know when you'd like a hint (if you'd like one).

OpenStudy (anonymous):

Think of what you can do with \(\mathbb{Q}\)

OpenStudy (anonymous):

How about you start by thinking of a function from that interval to R, that is continuous nowhere.

OpenStudy (zarkon):

should i give it away

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I'm going to give a hint by defining a function continuous nowhere. One second...

OpenStudy (anonymous):

\[ f(x) = \left\{ \begin{array}{c|c} 1 & x \in \mathbb{Q}\\ 0 & x \notin \mathbb{Q} \end{array}\right. \]

OpenStudy (across):

\[f(x)=\frac{C}{x-x}\]This one is messed up everywhere.

OpenStudy (anonymous):

That's not defined anywhere. The problem states that the function must be defined on the interval.

OpenStudy (zarkon):

that is not defined on [0,1]

OpenStudy (across):

I know it's not... I'm just brainstorming. xd

OpenStudy (anonymous):

Start by having a look at my function above. You can play around with it to get the solution.

OpenStudy (zarkon):

modify alchemista's hint

OpenStudy (across):

What's the syntax for those brackets?

OpenStudy (anonymous):

The source for my function: f(x) = \left\{ \begin{array }{c|c} 1 & x \in \mathbb{Q}\\ 0 & x \notin \mathbb{Q} \end{array}\right.

OpenStudy (anonymous):

Remove the space in \begin{array }

OpenStudy (anonymous):

You can right click the source for anyone's latex to show the source and copy it fyi

OpenStudy (across):

sec... syntax errors...

OpenStudy (anonymous):

If you've thought of a solution, just paste the source and I'll correct the errors in the LaTeX

OpenStudy (anonymous):

When you are thinking about this, remember about the delta-epsilon definition of continuity. \(f(x)\) is continuous at \(a\) if for every \(\varepsilon >0\) there is a \(\delta > 0\) such that when \(|x - a| < \delta\) then \(|f(x) - f(a)| < \varepsilon\)

OpenStudy (across):

\[ f(x) =\left \{ \begin {array} {c|c} 1 & x \in [0.5, 0.5] \\ 0 & x \notin (-\infty, 0) \cup (1, \infty) \end {array} \right \]

OpenStudy (across):

ich gebe auf xd

OpenStudy (anonymous):

\[f(x) =\left \{ \begin {array} {c|c} 1 & x \in [0.5, 0.5] \\ 0 & x \notin (-\infty, 0) \cup (1, \infty) \end {array} \right.\]

OpenStudy (anonymous):

This?

OpenStudy (anonymous):

The problem with your function is, it is only defined at 0.5 on the required interval.

OpenStudy (anonymous):

The function must be defined everywhere on [0,1] but must be continuous only at one point.